OpenStudy (anonymous):

Help please! Write 135% as a decimal. Write 0.5% as a decimal.

hero (hero):

\[\frac{135}{100} = \frac{100 + 35}{100} = \frac{100}{100} + \frac{35}{100} = 1 + \frac{35}{100} = 1.35\]

OpenStudy (anonymous):

Thanks!

hero (hero):

\[\frac{.05}{100} = \frac{5/100}{100} = \frac{5}{100} \times \frac{1}{100} = .05 \times .01 = 0.0005\]